OpenStudy (anonymous):

Derivative of f(a) a^4 sin2a tan2a ? So far I know 3a^3, but confused on how to proceed for sin/tan.

OpenStudy (anonymous):

4a^3*

OpenStudy (anonymous):

\[f(a)= a ^{4} \sin2a \tan2a\]

OpenStudy (calculusfunctions):

\[f \prime = 4a ^{3}\sin 2a \tan 2a + a ^{4}(2\cos 2a)(\tan 2a) + a ^{4}(\sin 2a)(2\sec ^{2}2a)\] You can simplify further. Was this helpful?

OpenStudy (anonymous):

So first I change the double angle formulas to their identities then I can do each derivative?

OpenStudy (anonymous):

 differentiate using product rule